OpenStudy (anonymous):

The sum of two numbers is 57. One of the numbers is 21. Which of the following equations could be used to find the other number?

OpenStudy (anonymous):

x + y = 57 x = 21 find y

jhonyy9 (jhonyy9):

57=21+x

OpenStudy (hoblos):

x=57 -21

OpenStudy (anonymous):

explain more hob

OpenStudy (anonymous):

x = 36

OpenStudy (hoblos):

the sum of two numbers x and y is 57 => x+y=57 if y=21 => x+21 = 57 => x=57-21

OpenStudy (amistre64):

\[n_1+n_2=57\]